OpenStudy (anonymous):

Simplify the expression. Assume that the variable represents a positive number square root 50x^4

OpenStudy (vijay):

5\[5\sqrt{2} X ^{2}\]

OpenStudy (anonymous):

Separate it into two groups with 2 diff. sqrt signs- perfect squares and non perfect squares \[\sqrt{25x^4}\sqrt{2}\]

OpenStudy (anonymous):

Take the sqrt of the perfect square root ones. \[5x^2\sqrt{2}\]

OpenStudy (anonymous):

done.